OpenStudy (anonymous):

Find the equation in point slope form of the line that is perpendicular to the line y=2x-3 and passes through the point (1,2)

OpenStudy (anonymous):

First, do you know the formula to use

OpenStudy (anonymous):

the point slope formula is y-y1=m(x-x1)

OpenStudy (calculusxy):

Point slope form: \(y - y_1 = m(x - x_1)\) perpendicular lines have an opposite (negative) reciprocal. e.g. \(2 \rightarrow -\frac{ 1 }{ 2 }\) \(\color{green}{\text{One-half is the perpendicular line's slope.}}\) Find the new y-intercept by plugging in the x and y values from the coordinate \((1,2)\). \(2 = m(1) + b\) and the \(m\) will be the perpendicular line's slope.

OpenStudy (anonymous):

I do not know the formula @Abbs___

OpenStudy (calculusxy):

The formula was stated by both of us: Once again it's: \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

the formula is y-y1=m(x-x1)

OpenStudy (anonymous):

Oops sorry but anyways thanks both of you. @calculusxy and @Abbys___